OpenStudy (davidusa):

Two of the three missing digits of the five-digit perfect square number 3_ ,__ 9 are the same. What is the other missing digit? MEDAL

OpenStudy (davidusa):

This question is from MathCounts.

OpenStudy (davidusa):

COME ON SOMEONE ANSWER MY QUESTION. THE ANSWER IS 0.

OpenStudy (raden):

117^2 = 31329

OpenStudy (davidusa):

2 of the numbers need to be the same. READ.

OpenStudy (raden):

try again :) 197^2 = 38809

OpenStudy (davidusa):

yes thats right how to get it?

OpenStudy (raden):

just trials

OpenStudy (davidusa):

This is mathcounts You have to do it fast. I need to know how to do it. How do you get it? with 2 digits the same?

OpenStudy (raden):

first, i use 200^2 get 40,000. ths closed but out of range look the unit number is 9, so my guess the units must be 3 or 7 finally, i take 197

OpenStudy (davidusa):

i still dont understand.

OpenStudy (raden):

@whpalmer4 , @mathmale , @ganeshie8 they are the good explainers

OpenStudy (davidusa):

ok thanks

OpenStudy (davidusa):

hello whpalmer4 thanks

OpenStudy (whpalmer4):

I think @RadEn covered it pretty well. 200^2 = 40000, so that's about spot where you need to start looking. If it is a perfect square ending in 9, you're only going to get that if the number being squared ends in a digit where its square ends in 9. 1*1 = 1 no good 2*2 = 4 no good 3*3 = 9 good 4*4 = 16 no good 5*5 = 25 no good 6*6 = 36 no good 7*7 = 49 good 8*8 = 64 no good 9*9 = 81 no good 0*0 = 0 no good So, you might reasonably guess that the nearest numbers to 200 (from below) ending in 3 or 7 are your obvious candidates, which would be 193 and 197. You can speed up the computation a bit, perhaps, if you remember that \[(a-b)(a-b) = a^2-2ab+b^2\]Let \(a=200\) and you have \(40000-800b+b^2\) and now you just plug in \(b=3\) or \(b=7\) to test \(197\) or \(193\) respectively.

OpenStudy (davidusa):

can you guys please help me on my problem?

OpenStudy (whpalmer4):

And looking at the \(40000-800b\) part you can immediately rule out \(b=7\) because the value will be \(<35000\).

OpenStudy (whpalmer4):

@DavidUsa is there another problem I'm not seeing here?

OpenStudy (whpalmer4):

Actually, I didn't look carefully enough at the original problem, you can't immediately rule out \(b=7\) because we don't know the second digit.

OpenStudy (phi):

the only way I see how to do it is by trial and error. as people have pointed out the upper bound is 200 the lower bound is sqr(3)*100 or about 174 the unit digit must be either a 3 or 7. we have the following candidates: 177 183 187 193 197 not hard to check with a calculator.

OpenStudy (phi):

only 197^2 gives a repeated number : 38809 the "unrepeated" number is 0

OpenStudy (whpalmer4):

Doing it by hand, I'd start from the largest candidates first, simply because they are a bit quicker to calculate with my suggestion. Just a happy accident that you encounter the solution first!

OpenStudy (phi):

or being "rainman" could be helpful

OpenStudy (whpalmer4):

There's an entertaining Feynman story about mental calculation of cube roots here: http://www.ee.ryerson.ca/~elf/abacus/feynman.html